fr3f3rf3rf3f3f3f:

Find the standard deviation of the data 3,9,10,5,8,5,9,11,10,3,10

cpl106:

3

fr3f3rf3rf3f3f3f:

thank you

cpl106:

lmk if its correct

jhonyy9:

@cpl106 please explain how you get this answer thank you in advance

cpl106:

σ2 = Σ(xi - μ)2 N = (3 - 7.5454545454545)2 + ... + (10 - 7.5454545454545)2 11 = 88.727272727273 11 = 8.0661157024793 σ = √8.0661157024793 = 2.8400907912388

cpl106:

thats how i got it but its not the full thing bc i don't have all the proper signs up their and etc and i didn't type the entire problem out